If two lengths of a right triangle are 16 and 30, what could be the length of the third side? Explain.
Answer by lwsshak3(11628)

If two lengths of a right triangle are 16 and 30, what could be the length of the third side? Explain.
**
Assume the hypotenuse=30
The third side would b = √(30^2-16^2)=√(900-256)=√644≈25.34 (other leg)
..
Assume the other leg=30
The third side would be = √(30^2+16^2)=√(900+256)=√1156=34 (hypotenuse)